OpenStudy (anonymous):

Let (-3,2) be a point on the terminal side of θ . Find the exact values of sin θ, csc θ, and cot θ

OpenStudy (anonymous):

sin theta = opp/hyp csc theta = hyp/opp or reciprocal of sin cot theta = opp/adj to find the hyp - (-3)^2 + (2)^2 = c^2 9 + 4 = c^2 sqrt(13) = c sin = 2/sqrt(13) = 2sqrt(13)/13 csc = sqrt(13)/2 cot = -2/3